# LNPT: Label-free Network Pruning and Training

### v2

@@ III-B Learning Gap
Where $\boldsymbol{C}^{t}$ represents the coefficient matrix at time t. $\boldsymbol{C}^{t}$ aligns the gradient fit with the variations in $\boldsymbol{\ell}$, enabling them to move in harmony. In the training process with data fitting as the objective, the change in $\boldsymbol{d}$ can be expressed as:
@@
where $\eta$ represents the learning rate, Loss denotes the fitting loss, and $\nabla_{\boldsymbol{\theta}_{s}}$oss represents the fitting gradient. Therefore, concerning the change in $\boldsymbol{\ell}$:
@@
We aim for the transformation through $\boldsymbol{C}^{t}$ to make the variations in $\boldsymbol{C}^{t}\nabla_{\boldsymbol{\theta}_{s}}Loss$ stabilize during different training stages. Concerning the learning gap $\boldsymbol{\ell}$, if it steadily decreases as training iterations progress, the network will converge faster. This aligns with the observation in [4] that ”pruned networks learn faster and achieve higher test accuracy and better generalization than the original network.” Therefore, through the coefficient matrix $\boldsymbol{C}^{t}$, we provide an equivalent expression for the relationship between $\boldsymbol{\ell}$ and network generalization, and the presence of $\boldsymbol{C}^{t}$ supports the feasibility of pre-pruning.
However, solving the coefficient matrix for large-scale networks is computationally infeasible. An alternative method that can closely reflect the change $\Delta\boldsymbol{\ell}$ in $\boldsymbol{\ell}$ during the training process needs to be proposed. It’s worth noting that the fitting gradient is significantly affected by the individual features of batches, leading to increased network overfitting and significant fluctuations. Therefore, it often requires adding penalties to suppress the individual features of the fitting gradient [27], and regularization of the network is a primary means of improving generalization performance. This implies that constructing an expression for $\Delta\ell$ based on the fitting gradient is not reasonable.
@@ III-E Dynamism of Feature Maps
Recent works [33][34] have investigated the training dynamics of large deep neural networks under gradient descent. Let $\mathcal{X}$,$\mathcal{Y}$ represent the inputs and labels, respectively. Consider a neural network $f(\cdot;\boldsymbol{\theta})$ parameterized by weights $\boldsymbol{\theta}$ and let $f\left(\mathcal{X};\boldsymbol{\theta}\right)$ be the concatenated vector of predictions for all training inputs. We emphasize the time-dependence of $\boldsymbol{\theta}$, where the NTK at time t is defined as [35]:
@@
0:  pruning ratio $\mathcal{P}$,pruning mask $\mathcal{M}$,training data $\mathcal{D}$,output of teacher network $f\left(\boldsymbol{\theta}_{t}\right)$ and feature map $\boldsymbol{m}_{t}$,output of student network $f\left(\boldsymbol{\theta}_{s}\right)$,feature map of student $\boldsymbol{m}_{s}$,number of main training steps $N$,learning rate $\alpha$
1:  $\mathcal{D}_{b}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{b}~{}\mathcal{D}$
@@
4:  $\mathcal{M}=S\left(\boldsymbol{\theta}_{s}\right)<\tau$
5:  $\mathcal{L}_{total}=\mathcal{L}_{oh}+\alpha\mathcal{L}_{\boldsymbol{m}}$
6:  for $i$ from 1 to $N$ do
@@ IV-A Unstructured Pruning
Two network architectures, VGG and ResNet, were employed to analyze the performance of LNPT on image classification tasks under different pruning ratios. Specifically, the whole network with L2 regularization constraint was selected as the baseline, and the compression efficiency of LNPT single-shot pruning on VGG19 and ResNet32 networks was examined on datasets such as CIFAR10 and CIFAR100.
To control experimental variables as much as possible, the same data augmentation techniques (random cropping and random flipping) were applied to the training dataset, similar to other methods. We initialized convolutional layer parameters using Kaiming normal distribution, used the Momentum optimizer, started with an initial learning rate of 0.1, and applied cosine annealing for learning rate scheduling.Additionally, for calculating evaluation metrics, the same sampling approach was followed, which involved using ten times the number of training samples’ labels (randomly selecting ten samples for each label).
@@
The performance comparison of CIFAR datasets on VGG-19 and ResNet-32 is shown in Table I and Fig.4 . It is evident that LNPT outperforms SNIP and GraSP in most cases. SNIP gets closer to the baseline at lower compression rates. However, due to the significant difference between teacher and student feature maps at high compression rates, LNPT exhibits superior performance at high compression rates. For the CIFAR-10 dataset with fewer categories, LNPT achieves accuracies of 92.06% and 85.63% at 99% pruning rate for VGG-19 and ResNet-32, respectively. For the more complex CIFAR-100 dataset, LNPT achieves accuracies of 67.33% and 51.44% for VGG-19 and ResNet-32, respectively. Additionally, the compact structure of ResNet and the presence of residual connections make the differences in feature maps more pronounced, further highlighting the advantages of LNPT. It’s worth noting that the combination of one-hot loss and intermediate semantics is essential for training, and it has also been used in previous work [41][42].
@@
LNPT focuses on the gradient characteristics of feature maps, making it more sensitive to the tail-end weights and pruning more in those areas. However, at high pruning rates, the top layers of the network are smaller in size, and LNPT retains more weights in the top layers compared to GraSP. This suggests that the top weights are highly sensitive to feature maps, playing an essential role in feature extraction. Therefore, LNPT retains more top weights compared to GraSP, which, on the other hand, prunes more top weights, potentially creating bottlenecks that hinder the transmission of feature information.
@@ IV-B Structured Pruning
For the VGG-16 network, LNPT can reduce FLOPs by 67.9%, and the pruned network still maintains high performance.90% of the parameters were pruned, and the network’s performance remained at a high level. In the case of the ResNet-56 network, nearly half of the network’s parameters (45.2%) were pruned, resulting in an increase of 0.24% in the accuracy of the pruned network. When FLOPs are reduced by 64.1%, the pruned network achieves an accuracy of 91.13%, with the pruned network’s parameters being only 25.9%.
Table IV presents the pruning results for different network structures on the CIFAR-100 dataset. Compared to CIFAR-10, CIFAR-100 has more classes (100 classes), making it more challenging and likely to lead to decreased network performance. However, LNPT still demonstrates good performance. Taking VGG16 and ResNet-56 as examples, FLOPs are reduced by 37.5% and 43.2%, respectively, while accuracy decreases by only 0.1% and 0.27%, respectively.
@@ IV-D An empirical approach for $\boldsymbol{C}^{t}$
In III, we qualitatively measure the generalization performance of the pruned network by using $\Delta\boldsymbol{m}$ to represent $\boldsymbol{\ell}$. In this section, an ideal approximation for $\boldsymbol{C}^{t}$ is provided, which is a focus of our future work. At time t, the change in $\Delta\boldsymbol{m}$ for the fitting target $Loss\left(\boldsymbol{\theta}_{s}^{t}\right)$ is represented as follows:
@@
$\boldsymbol{C}^{t}$ can provide a reasonable explanation for training before pruning. Previously, we aimed to preserve parameters sensitive to $\Delta\boldsymbol{d}$. However, since $\Delta\boldsymbol{d}$ cannot accurately reflect the network’s generalization ability over the long training process, pruning methods based on $\Delta\boldsymbol{d}$ cannot identify which parameters are globally effective. In the case of $\Delta\boldsymbol{\ell}$, since $\boldsymbol{\ell}$ steadily decreases throughout the training process, by preserving parameters sensitive to $\Delta\boldsymbol{\ell}$, parameters that are effective for generalization on a global scale can be identified.

### v3

@@ III-B Learning Gap
[…]
@@
[…]
@@
[…]
@@ III-E Dynamism of Feature Maps
Recent works [33][34] have investigated the training dynamics of large deep neural networks under gradient descent. Let $\mathcal{X}$$\mathcal{Y}$ represent the inputs and labels, respectively. Consider a neural network $f(\cdot;\boldsymbol{\theta})$ parameterized by weights $\boldsymbol{\theta}$ and let $f\left(\mathcal{X};\boldsymbol{\theta}\right)$ be the concatenated vector of predictions for all training inputs. We emphasize the time-dependence of $\boldsymbol{\theta}$, where the NTK at time t is defined as [35]:
@@
0: pruning ratio $\mathcal{P}$, pruning mask $\mathcal{M}$,training data $\mathcal{D}$, output of teacher network $f\left(\boldsymbol{\theta}_{t}\right)$ and feature map $\boldsymbol{m}_{t}$, output of student network $f\left(\boldsymbol{\theta}_{s}\right)$, feature map of student $\boldsymbol{m}_{s}$, number of main training steps NN, learning rate $\alpha$
[…]
@@
[…]
6:for $i$ from 1 to NN do
@@ IV-A Unstructured Pruning
Two network architectures, VGG and ResNet, were used to assess LNPT’s performance on image classification tasks across various pruning ratios. The entire network with L2 regularization constraint served as the baseline. LNPT single-shot pruning efficiency was evaluated on VGG19 and ResNet32 networks using datasets such as CIFAR10 and CIFAR100.
To control experimental variables, the same data augmentation techniques (random cropping and flipping) were applied to the training dataset. Convolutional layer parameters were initialized using Kaiming normal distribution, and the Momentum optimizer with an initial learning rate of 0.1 was used, along with cosine annealing for learning rate scheduling. Additionally, the evaluation metrics calculation employed the same sampling approach, with ten samples randomly selected for each label, amounting to ten times the number of training samples’ labels.
@@
The performance comparison of CIFAR datasets on VGG-19 and ResNet-32 is shown in Table I and Fig. 4. LNPT outperforms SNIP and GraSP in most cases. SNIP approaches the baseline at lower compression rates but falls short due to significant differences between teacher and student feature maps at high compression rates, where LNPT excels. For CIFAR-10 with fewer categories, LNPT achieves accuracies of 92.06% and 85.63% at 99% pruning rate for VGG-19 and ResNet-32, respectively. For the more complex CIFAR-100 dataset, LNPT achieves accuracies of 67.33% and 51.44% for VGG-19 and ResNet-32, respectively. ResNet’s compact structure and residual connections accentuate differences in feature maps, further showcasing LNPT’s advantages. Notably, the combination of one-hot loss and intermediate semantics is crucial for training, as seen in previous work [41][42].
@@
LNPT focuses on gradient characteristics of feature maps, being more sensitive to tail-end weights and pruning more in those areas. However, at high pruning rates, the top layers of the network become smaller, and LNPT retains more weights in the top layers compared to GraSP. This indicates the significance of top weights in feature extraction, as they are highly sensitive to feature maps. Consequently, LNPT retains more top weights compared to GraSP, which prunes more top weights, potentially creating bottlenecks that impede feature information transmission.
@@ IV-B Structured Pruning
For VGG-16, LNPT reduces FLOPs by 67.9%, while maintaining high performance in the pruned network. 90% of parameters were pruned, yet network performance remained high. In the case of ResNet-56, almost half of parameters (45.2%) were pruned, resulting in a 0.24% accuracy increase. With a 64.1% FLOPs reduction, the pruned network achieves 91.13% accuracy, with pruned parameters comprising only 25.9%.
Table IV shows pruning results for various network structures on CIFAR-100 dataset. Compared to CIFAR-10, CIFAR-100 has more classes (100 classes), making it more challenging and likely to decrease network performance. However, LNPT still performs well. Taking VGG16 and ResNet-56 as examples, FLOPs reduce by 37.5% and 43.2%, respectively, with only a 0.1% and 0.27% decrease in accuracy.
@@ IV-D An empirical approach for $\boldsymbol{C}^{t}$
[…]
@@
[…]